OpenStudy (anonymous):

Good afternoon, please how can I integrate this function? dx/(4x^2-9)^2

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

square it and you will get an easy polynomial to integrate

OpenStudy (anonymous):

can you explain more please? I thought that I can solve by trig. sustitu

OpenStudy (anonymous):

You can proceed with either a trigo sub or partial fractions. Let's try the former. Set \(x=\dfrac{3}{2}\sec t\), then \(dx=\dfrac{3}{2}\sec t\tan t\,dt\). \[\begin{align*} I&=\int\frac{dx}{(4x^2-9)^2}\\\\ &=\frac{3}{2}{\Huge\int}\frac{\sec t\tan t}{\left(4\left(\frac{3}{2}\sec t\right)^2-9\right)^2}\,dt\\\\ &=\frac{3}{2}{\Large\int}\frac{\sec t\tan t}{\left(4\times\frac{9}{4}\sec^2 t-9\right)^2}\,dt\\\\ &=\frac{3}{2}{\Large\int}\frac{\sec t\tan t}{9^2\left(\sec^2 t-1\right)^2}\,dt\\\\ &=\frac{1}{54}{\Large\int}\frac{\sec t\tan t}{\left(\tan^2 t\right)^2}\,dt\\\\ &=\frac{1}{54}\int\frac{\sec t}{\tan^3 t}\,dt\\\\ &=\frac{1}{54}\int\frac{\cos^2t}{\sin^3t}\,dt\\\\ &=\frac{1}{54}\int\frac{1-\sin^2t}{\sin^3t}\,dt\\\\ &=\frac{1}{54}\left(\int\csc^3t\,dt-\int\csc t\,dt\right) \end{align*}\] You can integrate by parts to derive the reduction formula for the integral of \(\csc^nt\) below: \[\int\csc^n t\,dt=-\frac{\csc^{n-2}t\cot t}{n-1}+\frac{n-2}{n-1}\int\csc^{n-2}t\,dt\] This gives \[\begin{align*} I&=\frac{1}{54}\left(-\frac{\csc t\cot t}{2}+\frac{1}{2}\int\csc t\,dt-\int\csc t\,dt\right)\\\\ &=-\frac{1}{108}\left(\csc t\cot t+\int\csc t\,dt\right) \end{align*}\]

OpenStudy (anonymous):

The latter approach, using partial fractions: \[\begin{align*} \frac{1}{(4x^2-9)^2}&=\frac{1}{(2x-3)^2(2x+3)^2}\\\\ &=\frac{A}{2x-3}+\frac{B}{(2x-3)^2}+\frac{C}{2x+3}+\frac{D}{(2x+3)^2}\\\\ 1&=A(2x-3)(2x+3)^2+B(2x+3)^2\\ &\quad\quad+C(2x-3)^2(2x+3)+D(2x-3)^2\\\\ &=A(8 x^3+12 x^2-18 x-27)+B(4 x^2+12 x+9)\\ &\quad\quad+C(8 x^3-12 x^2-18 x+27)+D(4 x^2-12 x+9)\\\\ &=(8A+8C)x^3+(12A+4B-12C+4D)x^2\\ &\quad\quad+(-18A+12B-18C-12D)x+(-27A+9B+27C+9D) \end{align*}\] Matching coefficients yields the system below, with solution: \[\begin{cases}8A+8C=0\\ 12A+4B-12C+4D=0\\ -18A+12B-18C-12D=0\\ -27A+9B+27C+9D=1\end{cases}~~\implies~~\begin{matrix}A=-\dfrac{1}{108}&&& B=\dfrac{1}{36}\\ C=\dfrac{1}{108}&&& D=\dfrac{1}{36}\end{matrix}\] To summarize, \[\begin{align*}\frac{1}{(4x^2-9)^2}&=-\frac{1}{108}\left(\frac{1}{2x-3}-\frac{3}{(2x-3)^2}-\frac{1}{2x+3}-\frac{3}{(2x+3)^2}\right)\\\\ \int\frac{1}{(4x^2-9)^2}\,dx&=-\frac{1}{108}\left(\int+\cdots+\int\right)\end{align*}\]